Determine the​ upper-tail critical value t α/2 in each of the following circumstances.

A. 1-a=0.95,n=27

B.1-a=0.99,n=27

C.1-a=0.95,n=16

D.1-a=0.95,n=41

E. 1-a=0.90,n=65
Answer #1

a) Lao 5.27 t0. 025.27-2.05 183

b) t_{\frac{1-0.99}{2},27} = t_{0.005,27} = 2.77068

c) t_{\frac{1-0.95}{2},27} = t_{0.025,16} = 2.11991

d) t_{\frac{1-0.95}{2},41} = t_{0.025,41} = 2.01954

e) t_{\frac{1-0.9}{2},65} = t_{0.05,65} = 1.66864